A patient is currently taking 220 mg of anhydrous zinc sulfate. To receive the equivalent amount of elemental zinc, how many milligrams of zinc sulfate heptahydrate (•7 H2O) would the patient have to take? (Molecular weights: zinc 65, ZnSO4 161, H2O 18
1
Expert's answer
2016-09-12T06:16:04-0400
w(Zn)ZnSO4 = M(Zn)/M(ZnSO4) * 100% = 65/161 * 100% = 40.3 %
w(Zn)ZnSO4*7H2O = M(Zn)/M(ZnSO4*7H2O) * 100% = 65/(161+18*7) * 100% = 22.6 %
m(Zn) = m(ZnSO4) * w(Zn)ZnSO4 = 220 * 0.403 = 88.66 (mg)
m (ZnSO4*7H2O) = m(Zn) / w(Zn)ZnSO4*7H2O = 88.66 / 0.226 = 392.3 (mg)